Evaluate the given expression and express the result using the usual format for writing numbers (instead of scientific notation)
41C2
this is what i tried 45!/(45!-2)but I think its wrong

Answer by stanbon(75887)

Evaluate the given expression and express the result using the usual format for writing numbers (instead of scientific notation)
41C2 = 41!/[(41-2)!*2!] = (41*40)/(1*2) = 41*20 = 820
==========================
Cheers,
Stan H.